OpenStudy (anonymous):

Two cars leave Denver at the same time and travel in opposite directions. one car travels 5 mi/h faster then the other car. the cars are 570 mi apart in 6h. how fast is each car traveling?

OpenStudy (anonymous):

570/6=95m/h 95/2=47.5 47.5-5=42.5miles/hr 47.5+5=52.5miles/hr

OpenStudy (anonymous):

omg i screwed up again

OpenStudy (anonymous):

45 and 50 i believe

OpenStudy (mertsj):

Do you know that rate multiplied by time equals distance?

OpenStudy (anonymous):

plz disregard what i did with the dividing up there it is suppose to be 95-5=90 90/2=45miles/hr 90-45=50 miles/hr

OpenStudy (radar):

Let x = speed of one car then x + 5 = speed of faster car. x + x + 5 = speed of one car relative to the other (they are traveling in opposite direction. Or 2x+5 Now d/s = t 570/(2x + 5) = 6 solving for x: 570= 12x + 30 12x = 540 x= 45 mph x+5 = 50 mph